OpenStudy (anonymous):

17/8 = e^(ln7/8 t) here, don't e and ln cancel out each other?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

can u try solving for t?

OpenStudy (anonymous):

because when i cancel them out, i got t= 2.42, which is different from the answer

OpenStudy (anonymous):

ans given is -5.64

OpenStudy (anonymous):

i would not call that "cancellation" but yes, \(e^{\ln(x)}=x\)

zepdrix (zepdrix):

Is the contents of the log \(\dfrac{7}{8t}\) or \(\dfrac{7}{8}t\)?

OpenStudy (anonymous):

the 2nd one

zepdrix (zepdrix):

Hmm then your solution sounds correct... that's strange.. thinking :3

OpenStudy (anonymous):

the solution apples an extra ln

zepdrix (zepdrix):

OH I SEE! The t is NOT inside the log, that's the issue here.

zepdrix (zepdrix):

So you can't "cancel" out the e and log, because the t is in the way.

zepdrix (zepdrix):

\[\huge e^{t \cdot \ln\frac{7}{8}}\]See the problem?

OpenStudy (anonymous):

oh, so u can separate it as e^t + e^ln7/8 -> then here u can change to 7/8?

zepdrix (zepdrix):

Yah that's another way to approach it! :)

OpenStudy (anonymous):

and still need to apply ln to isolate t

OpenStudy (anonymous):

hah thanks zepdirx :D